OpenStudy (anonymous):

Find the indicated limit, if it exists. limit of f of x as x approaches negative 8 where f of x equals x plus 9 when x is less than negative 8 and f of x equals negative seven minus x when x is greater than or equal to negative 8 -8 1 17 The limit does not exist.

OpenStudy (anonymous):

i have no idea how to do this

OpenStudy (solomonzelman):

let me interpret the question: Find \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~-8}f(x)}\) ? where, \(\LARGE\color{black}{ f(x) = \begin{cases} & x+9,~~~~~~~~~~{\large x<-8} \\ & -7-x,~~~~~~{\large x\ge-8} \end{cases} }\)

OpenStudy (solomonzelman):

is this interpretation correct?

OpenStudy (anonymous):

correct

OpenStudy (solomonzelman):

so to find the left sided limit, you need to plug in -8, into x+9 and to find the right sided limit, you need to plug in -8, into -7-x.

OpenStudy (solomonzelman):

if the two sides of the limit are no equivalent, then \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~-8}f(x)~~DNE}\)

OpenStudy (anonymous):

it equal;s 1

OpenStudy (solomonzelman):

yup!

OpenStudy (anonymous):

oh. so to find the limit, just plug in the numbers that they give you

OpenStudy (solomonzelman):

basically, but you need to verify that \(\Large\color{blue}{\displaystyle\lim_{x \rightarrow ~-8^{-}}f(x)=\displaystyle\lim_{x \rightarrow ~-8^{+}}f(x)}\)

OpenStudy (anonymous):

oh. right.

OpenStudy (solomonzelman):

if, \(\Large\color{blue}{\displaystyle\lim_{x \rightarrow ~-8^{-}}f(x) \ne\displaystyle\lim_{x \rightarrow ~-8^{+}}f(x)}\) then \(\Large\color{blue}{\displaystyle\lim_{x \rightarrow ~-8}f(x)~~\color{black}{\rm DNE}}\)

OpenStudy (anonymous):

i have a couple of more. do you mind?

OpenStudy (solomonzelman):

in your case, as you have already said both parts are equal to 1:) so 1 is the answer you need

OpenStudy (solomonzelman):

yes, sure.....

OpenStudy (solomonzelman):

you can draw th piece wise functions and the limits, that would be faster for both of us:)

OpenStudy (anonymous):